Find the x- and y- intercepts for the graph of the following equation.
5x + 6y = 120

Answer by nerdybill(7384)

Find the x- and y- intercepts for the graph of the following equation.
5x + 6y = 120
.
x-intercept is found by setting y to 0 and solving for x:
5x + 6y = 120
5x + 6(0) = 120
5x = 120
x = 24
x-intercept is at (24,0)
.
y-intercept is found by setting x to 0 and solving for y:
5x + 6y = 120
5(0) + 6y = 120
6y = 120
y = 20
y-intercept is at (0, 20)